# SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation of Large Language Models

### v1

@@ Abstract
The rapid advancement in large language models (LLMs) comes with a significant increase in their parameter size, presenting challenges for adaptation and fine-tuning. Parameter-efficient fine-tuning (PEFT) methods are widely used to adapt LLMs for downstream tasks efficiently. In this paper, we propose Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA, a novel PEFT method. We introduce a method to analyze the variation of the parameters by performing singular value decomposition (SVD) on weights and discuss SORSA’s superiority in minimizing the deviation from the pre-trained weight. Each SORSA layer consists of two main parts: trainable principle singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$. These parts are initialized by performing SVD on pre-trained weights. Moreover, we implement an orthonormal regularizer and analyze its importance by performing gradient analysis. The analysis shows that the regularizer could effectively transfer the scaling information into $\Sigma_{p}$, which ensures the parameter updating of SORSA layers is evenly and minimized on $U_{p}$ and $V^{\top}_{p}$. SORSA layers could be merged during inference, thus eliminating inference latency. After all, SORSA shows a faster convergence speed than PiSSA and LoRA in our experiments. On the MATH benchmark, Llama 2 7B adapted using SORSA achieved 10.36% accuracy, outperforming LoRA (5.50%), Full FT (7.22%), and PiSSA (7.44%). On the GSM-8K benchmark, SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%) We conclude that SORSA offers a new perspective on parameter-efficient fine-tuning, demonstrating remarkable performance. The code is available at https://github.com/Gunale0926/SORSA.
@@
Figure 1: Architecture of a SORSA layer. We only train parts rendered in orange ($U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$), and freeze parts rendered in blue ($U_{r}$, $\Sigma_{r}$ and $V^{\top}_{r}$).
@@ 1 Introduction
In this paper, we propose a novel PEFT approach, Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA. A SORSA layer has two main parts: principle singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$. These two parts are initialized by performing singular value decomposition (SVD) on pre-trained weight. Residual singular values and vectors will be merged into one matrix and frozen while training. We only train principle singular values and vectors with an orthonormal regularizer implemented to keep the orthonormality of $U_{p}$ and $V_{p}^{\top}$. The architecture of a SORSA layer is illustrated in Figure 1.
Furthermore, we analyze the pattern of variance of singular values and vectors during parameter updating and discuss the different patterns of partial fine-tuning (FT), LoRA, SORSA without regularizer, and SORSA with regularizer concerning singular values and vectors’ updating.
@@ 2 Related Works
LoRA (Hu et al., 2021) and its variants are the most popular type of PEFT methods. This type of PEFT is popular for its on-par or better performance than full parameter fine-tuning without introducing any inference latency. LoRA could be represented by equation $W=W_{0}+BA$, where $W_{0}\in\mathbb{R}^{m\times n}$ is the pre-trained weight, $A\in\mathbb{R}^{m\times r}$, using Gaussian initialization, and $B\in\mathbb{R}^{r\times n}$, using zero initialization, are low-rank matrices. Its variant, for example, AdaLoRA (Zhang et al., 2023) introduces an SVD decomposition and pruning for least significant singular values for more efficient parameter updating; DoRA (Liu et al., 2024) proposed a novel way to decompose weight into direction and magnitude by $W=\underline{m}\frac{W_{0}+\underline{BA}}{\|W_{0}+\underline{BA}\|_{c}}$, where $\underline{m}$ is initialized by $\underline{m}=\|W_{0}+\underline{BA}\|_{c}$, $\|\cdot\|_{c}$ denotes column-wise norm. The results show that DoRA has a better learning capacity than LoRA. However, DoRA introduced a calculation of norm in every training step, which makes it much more inefficient in training than LoRA; PiSSA (Meng et al., 2024) decomposes pre-trained weight $W_{0}$ into frozen $W_{pri}=AB$ and $W_{res}=W_{0}-W_{pri}$ based on singular values, and only updating the most important part $W_{pri}$, which result a faster convergence speed and better fitting than LoRA. SORSA has a similar architecture to PiSSA, which both conducts singular value decomposition (SVD) and replaces pre-trained weights with residual singular weights. SORSA inherits LoRA and its variants’ benefits, including low training VRAM requirement, no inference burden, and versatility in different architectures.
Other methods. There are also a few PEFT methods with unique techniques. For example, GaLore (Zhao et al., 2024) is a memory-efficient PEFT method that reduces VRAM usage by leveraging gradient accumulation and low-rank approximation. Despite its efficiency, GaLore’s complex implementation, involving Singular Value Decomposition (SVD) on gradients, adds computational complexity and may face scalability issues for very large models. LISA (Pan et al., 2024) uses a layer-wise importance sampling approach, prioritizing layers that significantly impact model performance. However, LISA’s selective fine-tuning requires careful resource allocation and introduces additional training complexity, complicating the pipeline and necessitating more sophisticated monitoring mechanisms.
@@ 3.2 Analysis Method
The study of DoRA (Liu et al., 2024) introduces an analysis method that focuses on the deviation of magnitude and direction ($\Delta M,\Delta D$) during training of full parameter fine-tuning and LoRA (Hu et al., 2021). They discovered that the distinction between full parameter fine-tuning and LoRA likely affects their learning ability difference. Inspired by their methods, we propose a novel method that analyzes the correlation between the deviation of singular values ($\Delta\Sigma$) and singular vectors ($\Delta D$) from pre-trained matrices. Our analysis suggests a significant difference in singular values and vectors’ stability and an updating pattern of partial fine-tuning, LoRA, and SORSA.
@@
$\Delta U_{j}=\left|\sum_{i=1}^{m}\left(U_{t}\odot U_{0}\right)_{ij}\right|$ (2)
$\Delta V^{\top}_{i}=\left|\sum_{j=1}^{n}\left(V_{t}^{\top}\odot V_{0}^{\top}\right)_{ij}\right|$ (3)
$\Delta D_{t}=1-\frac{1}{2k}\sum_{i=0}^{k}(\Delta U_{i}+\Delta V^{\top}_{i})$ (4)
Here, $\Delta D_{t}\in(0,1)$ represents variants of singular vectors between $W_{0}$ and $W_{t}$ at training step $t$. $\Delta U\in\mathbb{R}^{k}$, $\Delta V^{\top}\in\mathbb{R}^{k}$. $\odot$ denotes element-wise product, and $\left|\cdot\right|$ denotes element-wise absolute value. $U_{t}$ and $V_{t}$ are decomposed from $W_{t}$ by performing SVD, $k=\min(m,n)$.
We adopt the analysis on Llama 2 7B (Touvron et al., 2023) using the first 100K data of MetaMathQA (Yu et al., 2024). We test partial fine-tuning, LoRA, and SORSA (with regularizer with $\gamma=0.1$ and without regularizer). See Appendix A.1 for training details of the analysis.
@@ 3.3 Analysis Result
SORSA with regularizer shows $\Delta D$ and $\Delta\Sigma$ are in a much smaller range than other analysis methods. Moreover, different matrices demonstrated unrelated updating patterns, compared to all parallel-like other three methods, indicating that the updates in the SORSA are less constrained and potentially converge faster. In contrast, SORSA without the orthonormal regularizer exhibits a greater $\Delta D$ and $\Delta\Sigma$ alteration at each training step compared to SORSA with the regularizer. SORSA without regularizer also shows a linear-like updating pattern, similar to LoRA and partial FT. The result verifies the importance of regularizer in maintaining stability and minimizing the deviation from pre-trained matrices during the training process. These updating patterns indicate that SORSA maintains the characteristics of the pre-trained matrix better, thus potentially preserving the model’s generalization ability more effectively. This property allows SORSA layers to be trained with higher learning rates without showing noticeable over-fitting compared to other tuning methods.
@@ 4.1 Singular Value and Orthonormal Regularized Singular Vector Adaptation
Therefore, for a pre-trained weight $W_{0}\in\mathbb{R}^{m\times n}$, we could split it based on its singular value into principle weight $W_{p}$ and residual weight $W_{r}$, where $W_{p}$ contains the most important part of information of the matrix, and $W_{r}$ contains the least significant part:
@@
Therefore, parameter updating of $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ in a SORSA layer at training step $t$ could be expressed as:
@@ 4.2 Gradient Analysis
In this section, we analyze the impact of an orthonormal regularization on the updates of the singular vector matrix $U_{p}$ and $V^{\top}_{p}$ during training and the importance of keeping $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ separately. We also analyze the scaling factor for the learning rate of $\Sigma_{p}$
Given $U_{p,t-1}\in\mathbb{R}^{m\times r}$ is an orthonormal matrix, which corresponds to the initial state of $U_{p}$ due to the property of SVD. Here, we denote the gradient without regularizer for matrix $W$ at step $t$ is by $\nabla_{W_{t}}$, and the gradient with regularizer by $\nabla^{\prime}_{W_{t}}$:
$\nabla_{U_{p,t-1}}=\nabla_{U_{p,t-1}}\mathcal{L}_{train}$ (12)
$\nabla^{\prime}_{U_{p,t-1}}=\nabla_{U_{p,t-1}}\mathcal{L}_{train}+\gamma\nabla_{U_{p,t-1}}\mathcal{L}_{reg}$ (13)
Where $\nabla_{U_{p,t-1}}\mathcal{L}_{reg}=\frac{\partial\mathcal{L}_{reg}}{\partial U_{p,t-1}}=0$ due to the orthonormality of $U_{p,t-1}$; $\nabla_{U_{p,t-1}}\mathcal{L}_{train}=\frac{\partial\mathcal{L}_{train}}{\partial U_{p,t-1}}$.
Thus, we could calculate the weight at step $t$ without regularizer $U_{t}$, and with regularizer $U_{t}^{\prime}$:
$U_{p,t}^{\prime}=U_{p,t}=U_{p,t-1}-\eta\nabla_{U_{p,t-1}}\mathcal{L}_{train}$ (14)
Based on their weight on step $t$, we could calculate $W_{t+1}$ and $W^{\prime}_{t+1}$. Here, we only focus on the gradient of $U_{p}$ and $V^{\top}_{p}$, thus ignoring the updating of $\Sigma_{p,t}$ in this step.
$U_{t+1}$ $=U_{p,t}-\eta\nabla_{U_{p,t}}\mathcal{L}_{train}$ (15)
$V^{\top}_{t+1}$ $=V^{\top}_{p,t}-\eta\nabla_{V^{\top}_{p,t}}\mathcal{L}_{train}$ (16)
$W_{t+1}=U_{t+1}\Sigma_{p,t+1}V^{\top}_{t+1}$ (17)
$U^{\prime}_{t+1}$ $=U_{p,t}-\eta\nabla_{U_{p,t}}\mathcal{L}_{train}-\gamma\nabla_{U_{p,t}}\mathcal{L}_{reg}$ (18)
$V^{\top\prime}_{t+1}$ $=V^{\top}_{p,t}-\eta\nabla_{V^{\top}_{p,t}}\mathcal{L}_{train}-\gamma\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}$ (19)
$W^{\prime}_{t+1}=U^{\prime}_{t+1}\Sigma_{p,t+1}V^{\top\prime}_{t+1}$ (20)
Under optimal circumstances, $U^{\prime}_{p,t+1}$ will be an orthonormal matrix. Therefore, $U^{\prime}_{p,t+1}$ can be interpreted as pure rotation, which does not change the norm of the input matrix. By enforcing the orthonormal property of $U_{p}$ and $V^{\top}_{p}$ at each step, their regularized gradients $\nabla_{U_{p}}\mathcal{L}_{reg}$ and $\nabla_{V^{\top}_{p}}\mathcal{L}_{reg}$ will ensuring that the scaling information is not propagated into $U_{p,t}$ and $V^{\top}_{p,t}$. Consequently, only $\Sigma_{p}$ retains all scaling information.
To prove this statement, we could calculate the difference between $W^{\prime}_{t+1}$ and $W_{t+1}$:
$\begin{split}W^{\prime}_{t+1}-W_{t+1}&=-\gamma\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}U_{p,t}\Sigma_{p,t+1}\\ &\quad+\eta\gamma\nabla_{U_{p,t}}\mathcal{L}_{train}\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}\Sigma_{p,t+1}\\ &\quad-\gamma\nabla_{U_{p,t}}\mathcal{L}_{reg}\Sigma_{p,t+1}V^{\top}_{p,t}\\ &\quad+\eta\gamma\nabla_{U_{p,t}}\mathcal{L}_{reg}\nabla_{V^{\top}_{p,t}}\mathcal{L}_{train}\Sigma_{p,t+1}\\ &\quad+\gamma^{2}\nabla_{U_{p,t}}\mathcal{L}_{reg}\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}\Sigma_{p,t+1}\end{split}$ (21)
Here, we could find that each term has the $\Sigma_{p,t+1}$. This shows that by optimizing respected to regularizer loss at step $t+1$ and optimizing $\Sigma_{p,t+1}$ respected to train loss on step $t+2$, we could efficiently increase the orthonormality of $U_{p}$ and $V^{\top}_{p}$ while have negligible side-effect on convergence. In short, adding regularizer could transfer scaling information from $U_{p}$ and $V^{\top}_{p}$, which come from optimizing for $\mathcal{L}_{train}$, into $\Sigma_{p}$ allows the matrix scaling information updates more holistically, ultimately presenting a better generalization. In Figure 3, we illustrate the effect of $\gamma$ on orthonormality, training loss, and gradient norm. The figures present almost no difference in loss and gradient norm in different $\gamma$ values, which indicates the strong robustness of training of SORSA.
@@
Moreover, training along with an Adam-like (Kingma & Ba, 2015) optimizer could accelerate this process because its momentum on parameters of $U_{t}$ will become much smaller than momentum on parameters of $\Sigma_{t}$, which means the optimizer will halt the parameter updating towards the direction point to orthonormality increasing in $U_{p}$ and $V^{\top}_{p}$, and more likely to update the scaling information to $\Sigma_{p}$. By ensuring the scaling information only be updated into $\Sigma_{p}$, we could also ensure that the parameter distribution of $W_{p}$ is evenly distributed.

### v2

@@ Abstract
The rapid advancement in large language models (LLMs) comes with a significant increase in their parameter size, presenting challenges for adaptation and fine-tuning. Parameter-efficient fine-tuning (PEFT) methods are widely used to adapt LLMs for downstream tasks efficiently. In this paper, we propose Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA, a novel PEFT method. We introduce a method to analyze the variation of the parameters by performing singular value decomposition (SVD) and discuss and analyze SORSA’s superiority in minimizing the alteration in the SVD aspect. Each SORSA adapter consists of two main parts: trainable principal singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$. These parts are initialized by performing SVD on pre-trained weights. Moreover, we implement and analyze an orthonormal regularizer, which could effectively transfer the scaling information into $\Sigma_{p}$ and ultimately allows the training process to be more efficient. SORSA adapters could be merged during inference, thus eliminating any inference latency. After all, SORSA shows a faster convergence than PiSSA and LoRA in our experiments. On the MATH benchmark, Llama 2 7B adapted using SORSA achieved 10.36% accuracy, outperforming LoRA (5.50%), Full FT (7.22%), and PiSSA (7.44%). On the GSM-8K benchmark, SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%). We conclude that SORSA offers a new perspective on parameter-efficient fine-tuning, demonstrating remarkable performance. The code is available at https://github.com/Gunale0926/SORSA.
@@
Figure 1: Architecture of a SORSA adapter. We only train parts rendered in orange ($U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$), and freeze parts rendered in blue ($U_{r}$, $\Sigma_{r}$ and $V^{\top}_{r}$).
@@ 1 Introduction
In this paper, we propose a novel PEFT approach, Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA. A SORSA adapter has two main parts: principal singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$. These two parts are initialized by performing singular value decomposition (SVD) on pre-trained weight. Residual singular values and vectors will be merged into one matrix and frozen while training. We only train principal singular values and vectors with an orthonormal regularizer implemented to keep the orthonormality of $U_{p}$ and $V_{p}^{\top}$. The architecture of a SORSA adapter is illustrated in Figure 1.
Furthermore, we analyze the pattern of variation of singular values and vectors during parameter updating and discuss the different patterns of partial fine-tuning (FT), LoRA, SORSA without regularizer, and SORSA with regularizer concerning singular values and vectors’ updating.
@@ 2 Related Works
LoRA (Hu et al., 2021) and its variants are the most popular type of PEFT methods. This type of PEFT is popular for its on-par or better performance than full parameter fine-tuning without introducing any inference latency. LoRA could be represented by equation $W=W_{0}+BA$, where $W_{0}\in\mathbb{R}^{m\times n}$ is the pre-trained weight, $A\in\mathbb{R}^{m\times r}$, using Gaussian initialization, and $B\in\mathbb{R}^{r\times n}$, using zero initialization, are low-rank matrices. Its variant, for example, AdaLoRA (Zhang et al., 2023) introduces an SVD decomposition and pruning for least significant singular values for more efficient parameter updating. DoRA (Liu et al., 2024) proposed a novel way to decompose weight into direction and magnitude by $W=\underline{m}\frac{W_{0}+\underline{BA}}{\|W_{0}+\underline{BA}\|_{c}}$, where $\underline{m}$ is initialized by $\underline{m}=\|W_{0}+\underline{BA}\|_{c}$, $\|\cdot\|_{c}$ denotes column-wise norm. The results show that DoRA has a better learning capacity than LoRA. However, DoRA introduced a calculation of norm in every training step, which makes it much more inefficient in training than LoRA. PiSSA (Meng et al., 2024) decomposes pre-trained weight $W_{0}=U\text{diag}(S)V^{\top}$ by Singular Value Decomposition (SVD), and then splits $W_{0}$ into $W_{pri}$ and $W_{res}$: $W_{pri}=BA$ which is trainable, where $A=U_{[:r]}\text{diag}(S^{\frac{1}{2}}_{[:r]})$ and $B=\text{diag}(S^{\frac{1}{2}}_{[:r]})V^{\top}_{[:r]}$; $W_{res}=U_{[:r,:r]}\text{diag}(S_{[:r]})V^{\top}_{[:r,:]}$ which is frozen. PiSSA results a faster convergence speed and better fitting compared to LoRA. SORSA has a similar architecture to PiSSA, which both conducts SVD and replaces pre-trained weights with residual singular weights. SORSA inherits LoRA and its variants’ benefits, including low training VRAM requirement, no inference burden, and versatility in different architectures.
Other methods. There are also a few efficient adapting methods with unique techniques. For example, GaLore (Zhao et al., 2024) is a memory-efficient PEFT method that reduces VRAM usage by leveraging gradient accumulation and low-rank approximation. Despite its efficiency, GaLore’s complex implementation, involving perform SVD on gradients, adds computational complexity and may face scalability issues for very large models. LISA (Pan et al., 2024) uses a layer-wise importance sampling approach, prioritizing layers that significantly impact model performance, and selectively fine-tune essential parameters.
@@ 3.2 Analysis Method
The study of DoRA (Liu et al., 2024) introduces an analysis method that focuses on the deviation of magnitude and direction ($\Delta M,\Delta D$) during training of full parameter fine-tuning and LoRA (Hu et al., 2021). They discovered that the distinction between full parameter fine-tuning and LoRA likely affects their learning ability difference. Inspired by their methods, we propose a novel method that analyzes the correlation between the deviation of singular values ($\Delta\Sigma$) and singular vectors ($\Delta D$) from pre-trained matrices during updating. Our analysis suggests a significant difference in singular values and vectors’ stability and an updating pattern of partial fine-tuning, LoRA, and SORSA.
@@
$\Delta U_{t,j}=\left|\langle\mathbf{u}_{t}^{j},\mathbf{u}_{0}^{j}\rangle\right|$ (2)
$\Delta V^{\top}_{t,i}=\left|\langle\mathbf{v}_{t}^{i},\mathbf{v}_{0}^{i}\rangle\right|$ (3)
$\Delta D_{t}=1-\frac{1}{2k}\sum_{i=0}^{k}(\Delta U_{t,i}+\Delta V^{\top}_{t,i})$ (4)
Here, $k=\min(m,n)$; $\mathbf{u}_{t}^{j}$ denotes the $j$-th column vector of matrix $U_{t}$, and $\mathbf{v}_{t}^{i}$ denotes the $i$-th row vector of matrix $V_{t}^{\top}$; $\Delta D_{t}\in(0,1)$ represents variation of singular vectors between $W_{0}$ and $W_{t}$ at training step $t$; $U_{t}$ and $V_{t}^{\top}$ are decomposed from $W_{t}$ by performing SVD.
We adopt the analysis on Llama 2 7B (Touvron et al., 2023) using the first 100K data of MetaMathQA (Yu et al., 2024). We test partial fine-tuning, LoRA, and SORSA (with regularizer and without regularizer). See Appendix A.1 for training details of the analysis.
@@ 3.3 Analysis Result
SORSA with regularizer shows $\Delta D$ and $\Delta\Sigma$ are in a much smaller range than other analysis methods. Moreover, different matrices demonstrated unrelated updating patterns, compared to all parallel-like other three methods, indicating that the updates in the SORSA are less constrained and potentially converge faster. In contrast, SORSA without the orthonormal regularizer exhibits a greater $\Delta D$ and $\Delta\Sigma$ alteration at each training step compared to SORSA with the regularizer. SORSA without regularizer also shows a linear-like updating pattern, similar to LoRA and partial FT. The result verifies the importance of regularizer in maintaining stability and minimizing the deviation from pre-trained matrices during the training process. These updating patterns indicate that SORSA maintains the characteristics of the pre-trained matrix better, thus potentially preserving the model’s generalization ability more effectively. This property allows SORSA adapters to be trained with higher learning rates without showing noticeable over-fitting compared to other tuning methods.
@@ 4.1 Singular Value and Orthonormal Regularized Singular Vector Adaptation
Therefore, for a pre-trained weight $W_{0}\in\mathbb{R}^{m\times n}$, we could split it based on its singular value into principal weight $W_{p}$ and residual weight $W_{r}$, where $W_{p}$ contains the most important part of information of the matrix, and $W_{r}$ contains the least significant part:
@@
Therefore, parameter updating of $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ in a SORSA adapter at training step $t$ could be expressed as:
@@ 4.2 Gradient Analysis
In this section, we provides a gradient analysis for the SORSA with the orthonormal regularizer, focusing on proving the regularizer’s convexity and find the upper bound of the hyperparameter $\gamma$. We also demonstrates how the regularizer stabilizes the training and facilitates its generalization.
To prove SORSA with regularizer is optimizable, we need to prove the convexity of the regularizer, and then use Lipschitz continuity to find the upper bound of $\gamma$.
In order to analyze the regularizer according to Eq. (9), we begin by noticing that the Frobenius norm $\|\cdot\|_{F}$ is a convex function. We then examine the terms $U^{\top}_{p}U_{p}-I$ and $V_{p}V^{\top}_{p}-I$, which are quadratic forms in $U_{p}$ and $V_{p}$ , respectively. To assess their convexity, we compute the Hessian matrix of each quadratic term with respect to $U_{p}$ and $V_{p}$. For the term $U^{\top}_{p}U_{p}-I$ , we first express it as a function of $U_{p}$. The Hessian of a quadratic form $U^{\top}_{p}U_{p}$ is given by $2I$, where $I$ is the identity matrix of appropriate dimensions. This is because the second-order derivative of $U^{\top}_{p}U_{p}$ with respect to $U_{p}$ is constant and positive semidefinite, ensuring convexity. The second term could use the exact same process, which is also convex. Therefore, the regularizer is convex.
Then, we analyze the Lipschitz continuity of the gradients to establish an appropriate bound for $\gamma$. A function $f$ is Lipschitz continuous with constant $L$ if:
$\|\nabla f(W_{1})-\nabla f(W_{2})\|\leq L\|W_{1}-W_{2}\|$ (12)
According to Eq. (11), the gradient of combined loss function $\mathcal{L}(W_{p})$ is:
$\nabla_{W_{p,t}}\mathcal{L}=\nabla_{W_{p,t}}\mathcal{L}_{train}+\frac{\gamma}{\eta_{d}}\nabla_{W_{p,t}}\mathcal{L}_{reg}$ (13)
Assuming that $\nabla_{W_{p}}\mathcal{L}_{train}$ and $\nabla_{W_{p}}\mathcal{L}_{reg}$ are Lipschitz continuous with constants $L_{train}$ and $L_{reg}$, respectively, the Lipschitz constant for the combined loss function can be written as:
$L=L_{train}+\frac{\gamma}{\eta_{d}}L_{reg}$ (14)
When the learning rate is at maximum point, $\eta_{t}=\eta_{d}$, to ensure convergence of the algorithm, the learning rate $\eta_{d}$ must satisfy the following condition:
$\eta_{d}<\frac{2}{L_{train}+\frac{\gamma}{\eta_{d}}L_{reg}}$ (15)
Which could be written as:
$\gamma<\frac{2-\eta_{d}L_{train}}{L_{reg}}$ (16)
We can assume that the regularizer’s gradients scale with $\eta_{d}$ , meaning that larger updating step (due to a larger $\eta_{d}$) will lead to greater deviations from orthonormality, which increases $L_{reg}$. Conversely, smaller steps lead to a more gradual progression towards orthonormality, which reduces $L_{reg}$. Therefore, we could assume $L_{reg}\propto\eta_{d}$, and rewrite the inequality as:
$\gamma<\frac{2}{k\eta_{d}}-\frac{L_{train}}{k}$ (17)
This inequality provides an upper bound on $\gamma$ to ensure that the regularizer does not prevent convergence when $\eta_{t}$ is at its maximum value $\eta_{d}$. The inequality shows that $\gamma$ should have inversely proportional relationship to $\eta_{d}$ while selecting the hyperparameter in order to make sure the SORSA adapters could converge.
@@
In Figure 3, we could see that below the bound, $\gamma$ have almost no effect on loss curve, reflecting the robustness of SORSA.
The reinforcement of orthonormality of $U_{p}$ and $V^{\top}_{p}$ concentrates the scaling information into $\Sigma_{p}$, leads to a more holistic parameter updating process. The regularizer ensures that $U_{p}$ and $V^{\top}_{p}$ act as pure rotation matrices, while all magnitude updates are captured within $\Sigma_{p}$ by decoupling singular values from singular vectors,
Moreover, training along with an Adam-like (Kingma & Ba, 2015) optimizer could accelerate this process because its momentum on parameters of $U_{t}$ will become much smaller than momentum on parameters of $\Sigma_{t}$, which means the optimizer will halt the parameter updating towards the direction point to orthonormality increasing in $U_{p}$ and $V^{\top}_{p}$, and more likely to update the scaling information to $\Sigma_{p}$.
This separation stabilizes learning, as the singular vectors remain orthonormal, preventing distortions and over-fitting. Furthermore, concentrating the updates within $\Sigma_{p}$ allows for a more efficient gradient flow, facilitating smoother updates and ensuring that scaling changes reflect true data variation. Consequently, this structured update mechanism fosters better generalization and balanced learning throughout the training process.